# How to Train Your LLM Web Agent: A Statistical Diagnosis

### v3

@@ Abstract
11footnotetext: Equal contribution 1ServiceNow Research 2Mila–Quebec AI Institute 3Polytechnique Montréal 4HEC Montréal 5McGill University 6Univeristé de Montréal
Figure 1: Compute–performance frontier on MiniWoB++ (results averaged over two seeds). The blue curve shows pure SFT on teacher demonstrations. Warm-colored curves represent hybrid runs that branch off from SFT checkpoints and continue training with RL. Early transitions to RL push the Pareto frontier achieving higher success rates for the same compute and is the only approach able to achieve over $30\%$ improvement on both held-out goals (left) and held-out tasks (right) closing the gap between open and closed-source models. See fig.˜4 for the corresponding plot with Qwen2.5 7B.
@@ 1 Introduction
Most research centers on single-step tasks like code or math domains with rapid feedback and simplified credit assignment which fall short of real-world web environments requiring sequential decisions and long-horizon planning. Recent benchmarks like WebArena [29], WorkArena [10], OSWorld [26], and The Agent Company [27] have exposed how brittle current methods become under delayed rewards, sparse feedback, and compounding errors. Addressing these settings demands not just better agents, but reproducible, compute-efficient training pipelines an area we directly tackle.
However, building such pipelines is nontrivial. Modern LLM post-training often involves a combination of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), with performance sensitive to a large number of interacting hyperparameters [13]. Single-seed results are noisy and misleading, but exhaustive tuning and multi-seed sweeps (e.g. [2]) remain infeasible for most labs due to the cost of LLM training. This makes it all the more critical to design pipelines that are not only effective but statistically robust and accessible within realistic compute budgets.
@@
To study this compute-performance trade-off we use a two-stage training pipeline. First, a LLaMA 3.3 70B teacher model generates $K$ successful trajectories to warm-start a smaller LLaMA 3.1 8B student via SFT. We then branch of various SFT checkpoints where training continues with an on-policy RL phase using Group Relative Policy Optimization (GRPO) [9]. Our central objective is to determine the optimal compute allocation and hyperparameter mix for training web-based LLM agents. To this end, we run 1,370 training configurations, varying key hyperparameters and branching points between SFT and RL. We then apply bootstrap-based analysis to estimate the impact of different hyper-parameters on downstream performance and how they vary across branching SFT checkpoints. This data-driven approach enables us to identify important considerations to get the most out of each run. We use this method to show the optimal mix between SFT and RL in the MiniWob++ environment, achieving better task accuracy at a significantly lower cost. Additionally, we provide concrete recommendations on compute allocations between SFT and RL on the more demanding WorkArena environment.
@@
These findings are significant for two reasons. First, they give smaller research groups a reproducible, budget-aware playbook for pushing open LLM agents closer to state-of-the-art without scaling model size. Second, they address a slice of the broader reproducibility concerns recently highlighted in the RL community [3, 13], offering a template for rigorous, statistically-grounded hyperparameter tuning and compute allocation in LLM agent training.
@@ 2.2 Off‑policy boot‑strapping via SFT
SFT offers a high‑quality, low‑variance gradient but is inherently off‑policy which can lead to poor generalization [6].
@@ 2.4 On-Policy Improvement with Multi-Turn GRPO
After the initial SFT warmup phase, we continue training using on-policy RL using GRPO as the optimization algorithm. Like REINFORCE [21], GRPO maximizes the expected return using a policy gradient objective. However, GRPO introduces additional structure by leveraging per-goal advantage normalization and importance weighting, wherein each trajectory is associated with a specific goal in our context, referring to the seed of the target task. For a given goal $g$, the group-normalized advantage function is:
@@
where $\mathbf{R}_{t,g}=(r_{t,1},\dots,r_{t,G})$ is the set of rewards across the $G$ goals at timestamp $t$. In addition to this similar to proximal policy optimization (PPO) [2] an importance-ratio is often applied to the GRPO objective:
@@
Tokens with $A_{t,g}=0$ contribute no learning signal yet still consume memory. Dropping them yields a constant effective batch size and modestly accelerates training [28].
@@ 2.5 Curriculum through Variance-Aware Boltzmann Sampling
To promote steady learning progress, we design a curriculum that prioritizes challenging tasks, neither trivial nor too difficult [22]. Specifically, we select tasks according to a Boltzmann distribution centered around a target return $\mu_{\text{target}}$ which specifies the desired performance threshold, encouraging focus on partially mastered tasks, with a temperature parameter $\rho_{\text{Curr}}$ controlling the sharpness of the distribution, with lower values concentrating probability mass tightly around $\mu_{\text{target}}$.
@@ 3 Methodology
Our training pipeline consists of two sequential stages SFT followed by RL framed as a resource allocation problem. We also describe our hyperparameter sweep strategy and statistical analysis protocol that consolidates hundreds of runs into reliable conclusions. We evaluate our recipe along two axes: compute cost, measured in FLOPs (using the formula from [1]), and model performance, assessed on both unseen training goals and held-out testing tasks.
@@ 4 Experimental Setup
Our experiments focus on two benchmarks. The first is MiniWoB++, a suite of 30 medium-horizon web interaction tasks, where we observe that optimal policies typically complete tasks in 2 to 5 steps. The second is WorkArena [10], a more challenging benchmark of 33 enterprise knowledge-work tasks, where we observe optimal policies generally complete tasks in 3 to 10 steps. These benchmarks provide a representative spectrum of sequential decision-making challenges faced by interactive LLM agents. Both benchmarks are depicted in Figure˜7.
@@
MiniWoB++ provides raw HTML trees, whereas WorkArena leverages accessibility trees (AxTrees), which we truncate to 16k tokens to meet hardware constraints. The agent operates in a discrete action space composed of high-level UI primitives: noop, fill(node, text), click(node), type(node, text), select_option(node, option), scroll(node), and hover(node). This abstraction allows the agent to interact effectively with diverse web interfaces. All agents employ chain-of-thought prompting [25]. We also experiment with applying error log feedback, allowing the agent to receive explicit error messages when it takes invalid actions (during both training and testing).
@@
To manage the training pipeline, we use BrowserGym [8] for orchestrating Chromium-based web environments and structuring the agent’s action space, while AgentLab [8] handles agent design. Model fine-tuning is conducted with Torchtune [23], utilizing Fully Sharded Data Parallelism (FSDP) to enable scalable training across multiple GPUs. Given the high memory demands of long-sequence processing, we apply activation offloading and gradient checkpointing techniques, achieving approximately 40% reduction in memory usage.
@@ 5 Main Results and Compute Trade-Offs
Notably, this is the only strategy that closes the performance gap with closed-source models like GPT-4o. The trend is consistent across both held-out goals and held-out tasks: warm-started RL yields higher performance than either SFT or RL alone, reinforcing the importance of blending expert supervision with on-policy learning. These findings underscore the need to balance sample efficiency from demonstrations with the compute efficiency of on-policy learning, and the results generalize to Qwen2.5 7B (see fig.˜4).
@@ 6 Ablation and Sensitivity Analysis
Figure˜3 displays our findings, which we summarize as follows. Curriculum learning is beneficial when starting RL from scratch but becomes detrimental after warm-starting, likely because warm-started models already perform well on easy tasks, and curriculum forces them to overfocus on harder ones. Error log feedback helps when there’s no SFT but otherwise does not, likely because SFT warmup removes many common errors made by weaker models. A decoding temperature of 0.25 consistently yields the best results, striking a balance between exploration and exploitation; lower values led to under-exploration and were discarded. Grouped-relative advantage helps only after SFT, while using raw rewards works better when starting directly from the Instruct model, possibly due to how advantage scaling interacts with the initial weights. Zero-advantage filtering improves training across most settings by ensuring batches focus on informative updates. Standard-deviation normalized advantage, as noted by [17], seems to aid performance under less finetuned models and decrease in value the more finetuning is done. Importance ratio correction and trust region, though standard, also hurt models with little or no SFT, likely because conservative updates slow down learning. In contrast, for models that start from a stronger SFT checkpoint, these mechanisms can help stabilize training and avoid catastrophic updates. For the learning rate, the larger value (1e-6) generally worked better. Effective batch size of 512 appears to be generally be a safe and robust choice in all our experiments. Finally, regarding the discount rate, values above 0.9 work well for models with little or no SFT warmup, while heavily warm-started models benefit from a lower rate around 0.5 likely because it encourages the agent to optimize more aggressively on tasks it already handles well. A further hypothesis is that the large deviation in optimal hyperparameters between the base and warm-start models could be attributed to the significant entropy reduction introduced by SFT (see Appendix˜H).
@@ 7 Related Work
Building on the recognition of reproducibility challenges and unstable RL training of LLM agents, recent studies have proposed best practices for training LLM agents using RL methods. Dang and Ngo [7] recommend leveraging high quality data, balancing easy and hard problems, and controlling length generation with cosine reward. Yu et al. [28] promote higher clipping in the GRPO loss to promote diversity and avoid entropy collapse, dynamic sampling to improve training efficiency and stability, token level gradients for long CoT sequences, and overlong reward shaping to reduce reward noise.  Roux et al. [20] introduce tapered variant of importance sampling to speed up learning while maintaining stable learning dynamics. The proposed method (TOPR) allows the handling of both positive and negative examples in a fully offline setting. More generally, Hochlehnert et al. [13] emphasizes the need for greater methodological precision, particularly concerning decoding parameters, random seeds, prompt formatting, as well as the hardware and software frameworks, to guarantee transparent and thorough assessments of model performance. These practices are essential for developing robust and reproducible agents.
@@
Recent advancements have sought to bridge the gap in training LLM agents for multi-step environments, with approaches like WebRL [19] and SWEET-RL [30] demonstrating significant progress. WebRL employs a self-evolving curriculum to address the challenges of sparse feedback and task scarcity, successfully enhancing the performance of open LLMs in web-based tasks [19]. Similarly, SWEET-RL introduces a hierarchical structure that enables effective credit assignment over multiple turns, improving policy learning and generalization in collaborative reasoning tasks [30]. These studies collectively illustrate the necessity of adapting RL techniques to accommodate the complexities of multi-step interactions, paving the way for more capable and versatile LLM agents. Similar in spirit to our work, [4] propose an empiral study on the inference cost of trained LLM web agents. An extended version of the related work is provided in Appendix˜F.
@@
Recent work has begun closing the gap for LLM agents that must reason over multiple steps. WebRL [19] introduces a self-evolving curriculum that tackles sparse feedback and task scarcity, substantially boosting open-source agents on web tasks. SWEET-RL [30] adds a hierarchical credit-assignment scheme spanning many turns, improving both learning stability and generalization in collaborative reasoning. Together, these studies underscore the need to adapt RL techniques to the intricacies of long-horizon interaction, paving the way for more capable, versatile agents. Similar in spirit, [4] provide an empirical analysis of inference costs for trained LLM web agents. An extended discussion of related work appears in Appendix˜F.
@@ Appendix C Extended Learning and Saturation Analysis
  • Observation/action space mismatch: One of the important thing to note specifically in our web environment is the observation space which the agent uses is a bit different from the action space. Multiple times, the agent can see the correct action in the AxTree but the action space, the icon is not visible and to make it visible, the agent needs to scroll down and do the action. This mismatch causes huge problems [15]
  • UI Misuse: The agent tries to interact with items in the environment in ways that it is not designed. For example, the agent trying to fill in a checkbox with value True while it should just click on it. [18]
  • Repeating actions: A common issue we observed is the repetition of actions across multiple consecutive steps, often accompanied by verbose and redundant chains of thought. The agent frequently restates similar thoughts or re-executes the same actions unnecessarily, leading to inefficiencies and sometimes getting stuck in loops.  [18].
@@ Appendix D Deriving Compute Cost
We estimate FLOPs per token using the following formula, adapted from nvidia benchmarking[1]:
@@ Appendix E Benchmark Descriptions
Figure 7: Example tasks in MiniWoB++ [16] (top) and WorkArena [10] (bottom). MiniWoB consists of single-page simple tasks such as selecting a particular date and using a basic text editor, while WorkArena comprises multi-page complex tasks like filling forms and placing orders in an enterprise environment.
@@ Appendix F Extended Related Work
The reproducibility crisis in large language models (LLMs) and reinforcement learning (RL) has garnered increasing attention, particularly due to the reliance on single seed results that distort the perceived performance of models. The reproducibility challenge***https://reproml.org/ organized every year is a positive step towards addressing this. More concretely, Hochlehnert et al. [13] provide a critical examination of how such practices undermine the reliability of published findings, revealing that many reported gains are sensitive to implementation choices, such as random seeds and prompt formatting [13].
@@
Previous work in RL for LLMs has predominantly focused on single-step tasks, which have shown effectiveness in mathematical reasoning and code generation [28, 9, 20]. While these approaches exhibit promising results, they are limited in their applicability to real-world scenarios, which often require multistep decision-making capabilities. The narrow focus on bandit-style problems fails to address the complexities inherent in tasks that demand sequential interaction, highlighting a significant gap in the current research landscape.
@@
To assess the capabilities of LLM agents in more realistic environments, benchmarks such as WebArena [29], WorkArena [10, 5], the Agent Company [27], and OSWorld [26] have been designed to evaluate agents on multi-step tasks across various domains. These benchmarks expose the limitations of current LLM agents, revealing that while they may perform well in controlled settings, their performance in practical applications remains subpar, underscoring the need for further advancements in agent robustness and generalization to multi-step planning.
@@
Building on the recognition of reproducibility challenges and unstable RL training of LLM agents, recent studies have proposed best practices for training LLM agents using RL methods. Dang and Ngo [7] recommend leveraging high quality data, balancing easy and hard problems, and controlling length generation with cosine reward. Yu et al. [28] promote higher clipping in the GRPO loss to promote diversity and avoid entropy collapse, dynamic sampling to improve training efficiency and stability, token level gradients for long CoT sequences, and overlong reward shaping to reduce reward noise.  Roux et al. [20] introduce tapered variant of importance sampling to speed up learning while maintaining stable learning dynamics. The proposed method (TOPR) allows the handling of both positive and negative examples in a fully offline setting. More generally, Hochlehnert et al. [13] emphasizes the need for greater methodological precision, particularly concerning decoding parameters, random seeds, prompt formatting, as well as the hardware and software frameworks, to guarantee transparent and thorough assessments of model performance.
@@
Recent advancements have sought to bridge the gap in training LLM agents for multi-step environments, with approaches like WebRL [19] and SWEET-RL [30] demonstrating significant progress. WebRL employs a self-evolving curriculum to address the challenges of sparse feedback and task scarcity, successfully enhancing the performance of open LLMs in web-based tasks [19]. Similarly, SWEET-RL introduces a hierarchical structure that enables effective credit assignment over multiple turns, improving policy learning and generalization in collaborative reasoning tasks [30]. These studies collectively illustrate the necessity of adapting RL techniques to accommodate the complexities of multi-step interactions, paving the way for more capable and versatile LLM agents.

### v4

@@ Abstract
[…]
Figure 1: Compute–performance frontier on MiniWoB++ (results averaged over two seeds). The blue curve shows pure SFT on teacher demonstrations. Warm-colored curves represent hybrid runs that branch off from SFT checkpoints and continue training with RL. Early transitions to RL push the Pareto frontier achieving higher success rates for the same compute and is the only approach able to achieve over $30\%$ improvement on both held-out goals (left) and held-out tasks (right) closing the gap between open and closed-source models. See figure˜4 for the Qwen2.5 7B plot.
@@ 1 Introduction
Most research centers on single-step tasks like code or math domains with rapid feedback and simplified credit assignment which fall short of real-world web environments requiring sequential decisions and long-horizon planning. Recent benchmarks like WebArena (zhou2023webarena), WorkArena (workarena2024), OSWorld (OSWorld), and The Agent Company (xu2024theagentcompanybenchmarkingllmagents) have exposed how brittle current methods become under delayed rewards, sparse feedback, and compounding errors. Addressing these settings demands not just better agents, but reproducible, compute-efficient training pipelines an area we directly tackle.
However, building such pipelines is nontrivial. Modern LLM post-training often involves a combination of Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL), with performance sensitive to a large number of interacting hyperparameters (hochlehnert2025repro). Single-seed results are noisy and misleading, but exhaustive tuning and multi-seed sweeps (e.g. (abdin2025phi4reasoningtechnicalreport)) remain infeasible for most labs due to the cost of LLM training. This makes it all the more critical to design pipelines that are not only effective but statistically robust and accessible within realistic compute budgets.
@@
To study this compute-performance trade-off we use a two-stage training pipeline. First, a LLaMA 3.3 70B teacher model generates $K$ successful trajectories to warm-start a smaller LLaMA 3.1 8B student via SFT. We then branch of various SFT checkpoints where training continues with an on-policy RL phase using Group Relative Policy Optimization (GRPO) (deepseekr1). Our central objective is to determine the optimal compute allocation and hyperparameter mix for training web-based LLM agents. To this end, we run 1,370 training configurations, varying key hyperparameters and branching points between SFT and RL. We then apply bootstrap-based analysis to estimate the impact of different hyper-parameters on downstream performance and how they vary across branching SFT checkpoints. This data-driven approach enables us to identify important considerations to get the most out of each run. We use this method to show the optimal mix between SFT and RL in the MiniWob++ environment, achieving better task accuracy at a significantly lower cost. Additionally, we provide concrete recommendations on compute allocations between SFT and RL on the more demanding WorkArena environment.
@@
These findings are significant for two reasons. First, they give smaller research groups a reproducible, budget-aware playbook for pushing open LLM agents closer to state-of-the-art without scaling model size. Second, they address a slice of the broader reproducibility concerns recently highlighted in the RL community (agarwal2021deep; hochlehnert2025repro), offering a template for rigorous, statistically-grounded hyperparameter tuning and compute allocation in LLM agent training.
@@ 2.2 Off‑policy boot‑strapping via SFT
SFT offers a high‑quality, low‑variance gradient but is inherently off‑policy which can lead to poor generalization (chu2025sftmemorizesrlgeneralizes).
@@ 2.4 On-Policy Improvement with Multi-Turn GRPO
After the initial SFT warmup phase, we continue training using on-policy RL using GRPO as the optimization algorithm. Like REINFORCE (reinforce), GRPO maximizes the expected return using a policy gradient objective. However, GRPO introduces additional structure by leveraging per-goal advantage normalization and importance weighting, wherein each trajectory is associated with a specific goal in our context, referring to the seed of the target task. For a given goal $g$, the group-normalized advantage function is:
@@
where $\mathbf{R}_{t,g}=(r_{t,1},\dots,r_{t,G})$ is the set of rewards across the $G$ goals at timestamp $t$. In addition to this similar to proximal policy optimization (PPO) (abdin2025phi4reasoningtechnicalreport) an importance-ratio is often applied to the GRPO objective:
@@
Tokens with $A_{t,g}=0$ contribute no learning signal yet still consume memory. Dropping them yields a constant effective batch size and modestly accelerates training (yu2025dapo).
@@ 2.5 Curriculum through Variance-Aware Boltzmann Sampling
To promote steady learning progress, we design a curriculum that prioritizes challenging tasks, neither trivial nor too difficult (thakkar2023selfinfluence). Specifically, we select tasks according to a Boltzmann distribution centered around a target return $\mu_{\text{target}}$ which specifies the desired performance threshold, encouraging focus on partially mastered tasks, with a temperature parameter $\rho_{\text{Curr}}$ controlling the sharpness of the distribution, with lower values concentrating probability mass tightly around $\mu_{\text{target}}$.
@@ 3 Methodology
Our training pipeline consists of two sequential stages SFT followed by RL framed as a resource allocation problem. We also describe our hyperparameter sweep strategy and statistical analysis protocol that consolidates hundreds of runs into reliable conclusions. We evaluate our recipe along two axes: compute cost, measured in FLOPs (using the formula from (benchmarking)), and model performance, assessed on both unseen training goals and held-out testing tasks.
@@ 4 Experimental Setup
Our experiments focus on two benchmarks. The first is MiniWoB++, a suite of 30 medium-horizon web interaction tasks, where we observe that optimal policies typically complete tasks in 2 to 5 steps. The second is WorkArena (workarena2024), a more challenging benchmark of 33 enterprise knowledge-work tasks, where we observe optimal policies generally complete tasks in 3 to 10 steps. These benchmarks provide a representative spectrum of sequential decision-making challenges faced by interactive LLM agents. Both benchmarks are depicted in Figure˜7.
@@
MiniWoB++ provides raw HTML trees, whereas WorkArena leverages accessibility trees (AxTrees), which we truncate to 16k tokens to meet hardware constraints. The agent operates in a discrete action space composed of high-level UI primitives: noop, fill(node, text), click(node), type(node, text), select_option(node, option), scroll(node), and hover(node). This abstraction allows the agent to interact effectively with diverse web interfaces. All agents employ chain-of-thought prompting (wei2023chainofthoughtpromptingelicitsreasoning). We also experiment with applying error log feedback, allowing the agent to receive explicit error messages when it takes invalid actions (during both training and testing).
@@
To manage the training pipeline, we use BrowserGym (chezelles2025the) for orchestrating Chromium-based web environments and structuring the agent’s action space, while AgentLab (chezelles2025the) handles agent design. Model fine-tuning is conducted with Torchtune (torchtune), utilizing Fully Sharded Data Parallelism (FSDP) to enable scalable training across multiple GPUs. Given the high memory demands of long-sequence processing, we apply activation offloading and gradient checkpointing techniques, achieving approximately 40% reduction in memory usage.
@@ 5 Main Results and Compute Trade-Offs
Notably, this is the only strategy that closes the performance gap with closed-source models like GPT-4o. The trend is consistent across both held-out goals and held-out tasks: warm-started RL yields higher performance than either SFT or RL alone, reinforcing the importance of blending expert supervision with on-policy learning. These findings underscore the need to balance sample efficiency from demonstrations with the compute efficiency of on-policy learning, and the results generalize to Qwen2.5 7B (see figure˜4).
@@ 6 Ablation and Sensitivity Analysis
Figure˜3 displays our findings, which we summarize as follows. Curriculum learning is beneficial when starting RL from scratch but becomes detrimental after warm-starting, likely because warm-started models already perform well on easy tasks, and curriculum forces them to overfocus on harder ones. Error log feedback helps when there’s no SFT but otherwise does not, likely because SFT warmup removes many common errors made by weaker models. A decoding temperature of 0.25 consistently yields the best results, striking a balance between exploration and exploitation; lower values led to under-exploration and were discarded. Grouped-relative advantage helps only after SFT, while using raw rewards works better when starting directly from the Instruct model, possibly due to how advantage scaling interacts with the initial weights. Zero-advantage filtering improves training across most settings by ensuring batches focus on informative updates. Standard-deviation normalized advantage, as noted by (liu2025understandingr1zeroliketrainingcritical), seems to aid performance under less finetuned models and decrease in value the more finetuning is done. Importance ratio correction and trust region, though standard, also hurt models with little or no SFT, likely because conservative updates slow down learning. In contrast, for models that start from a stronger SFT checkpoint, these mechanisms can help stabilize training and avoid catastrophic updates. For the learning rate, the larger value (1e-6) generally worked better. Effective batch size of 512 appears to be generally be a safe and robust choice in all our experiments. Finally, regarding the discount rate, values above 0.9 work well for models with little or no SFT warmup, while heavily warm-started models benefit from a lower rate around 0.5 likely because it encourages the agent to optimize more aggressively on tasks it already handles well. A further hypothesis is that the large deviation in optimal hyperparameters between the base and warm-start models could be attributed to the significant entropy reduction introduced by SFT (see Appendix˜H).
@@ 7 Related Work
Building on the recognition of reproducibility challenges and unstable RL training of LLM agents, recent studies have proposed best practices for training LLM agents using RL methods. dang2025rlforreasoning recommend leveraging high quality data, balancing easy and hard problems, and controlling length generation with cosine reward. yu2025dapo promote higher clipping in the GRPO loss to promote diversity and avoid entropy collapse, dynamic sampling to improve training efficiency and stability, token level gradients for long CoT sequences, and overlong reward shaping to reduce reward noise. leroux2025tapered introduce tapered variant of importance sampling to speed up learning while maintaining stable learning dynamics. The proposed method (TOPR) allows the handling of both positive and negative examples in a fully offline setting. More generally, hochlehnert2025repro emphasizes the need for greater methodological precision, particularly concerning decoding parameters, random seeds, prompt formatting, as well as the hardware and software frameworks, to guarantee transparent and thorough assessments of model performance. These practices are essential for developing robust and reproducible agents.
@@
Recent advancements have sought to bridge the gap in training LLM agents for multi-step environments, with approaches like WebRL (qi2025webrl) and SWEET-RL (zhou2025sweetrltrainingmultiturnllm) demonstrating significant progress. WebRL employs a self-evolving curriculum to address the challenges of sparse feedback and task scarcity, successfully enhancing the performance of open LLMs in web-based tasks (qi2025webrl). Similarly, SWEET-RL introduces a hierarchical structure that enables effective credit assignment over multiple turns, improving policy learning and generalization in collaborative reasoning tasks (zhou2025sweetrltrainingmultiturnllm). These studies collectively illustrate the necessity of adapting RL techniques to accommodate the complexities of multi-step interactions, paving the way for more capable and versatile LLM agents. Similar in spirit to our work, (andreux2025surfer) propose an empiral study on the inference cost of trained LLM web agents. An extended version of the related work is provided in Appendix˜F.
@@
Recent work has begun closing the gap for LLM agents that must reason over multiple steps. WebRL (qi2025webrl) introduces a self-evolving curriculum that tackles sparse feedback and task scarcity, substantially boosting open-source agents on web tasks. SWEET-RL (zhou2025sweetrltrainingmultiturnllm) adds a hierarchical credit-assignment scheme spanning many turns, improving both learning stability and generalization in collaborative reasoning. Together, these studies underscore the need to adapt RL techniques to the intricacies of long-horizon interaction, paving the way for more capable, versatile agents. Similar in spirit, (andreux2025surfer) provide an empirical analysis of inference costs for trained LLM web agents. An extended discussion of related work appears in Appendix˜F.
@@ Appendix C Extended Learning and Saturation Analysis
  • Observation/action space mismatch: One of the important thing to note specifically in our web environment is the observation space which the agent uses is a bit different from the action space. Multiple times, the agent can see the correct action in the AxTree but the action space, the icon is not visible and to make it visible, the agent needs to scroll down and do the action. This mismatch causes huge problems VisualWebArena2024
  • UI Misuse: The agent tries to interact with items in the environment in ways that it is not designed. For example, the agent trying to fill in a checkbox with value True while it should just click on it. Murty2025NNetNav
  • Repeating actions: A common issue we observed is the repetition of actions across multiple consecutive steps, often accompanied by verbose and redundant chains of thought. The agent frequently restates similar thoughts or re-executes the same actions unnecessarily, leading to inefficiencies and sometimes getting stuck in loops. (Murty2025NNetNav).
@@ Appendix D Deriving Compute Cost
We estimate FLOPs per token using the following formula, adapted from nvidia benchmarkingbenchmarking:
@@ Appendix E Benchmark Descriptions
Figure 7: Example tasks in MiniWoB++ liu2018reinforcement (top) and WorkArena workarena2024 (bottom). MiniWoB consists of single-page simple tasks such as selecting a particular date and using a basic text editor, while WorkArena comprises multi-page complex tasks like filling forms and placing orders in an enterprise environment.
@@ Appendix F Extended Related Work
The reproducibility crisis in large language models (LLMs) and reinforcement learning (RL) has garnered increasing attention, particularly due to the reliance on single seed results that distort the perceived performance of models. The reproducibility challenge [^1] https://reproml.org/ organized every year is a positive step towards addressing this. More concretely, hochlehnert2025repro provide a critical examination of how such practices undermine the reliability of published findings, revealing that many reported gains are sensitive to implementation choices, such as random seeds and prompt formatting hochlehnert2025repro.
@@
Previous work in RL for LLMs has predominantly focused on single-step tasks, which have shown effectiveness in mathematical reasoning and code generation yu2025dapo; deepseekr1; leroux2025tapered. While these approaches exhibit promising results, they are limited in their applicability to real-world scenarios, which often require multistep decision-making capabilities. The narrow focus on bandit-style problems fails to address the complexities inherent in tasks that demand sequential interaction, highlighting a significant gap in the current research landscape.
@@
To assess the capabilities of LLM agents in more realistic environments, benchmarks such as WebArena zhou2023webarena, WorkArena workarena2024; boisvert2024workarenacompositionalplanningreasoningbased, the Agent Company xu2024theagentcompanybenchmarkingllmagents, and OSWorld OSWorld have been designed to evaluate agents on multi-step tasks across various domains. These benchmarks expose the limitations of current LLM agents, revealing that while they may perform well in controlled settings, their performance in practical applications remains subpar, underscoring the need for further advancements in agent robustness and generalization to multi-step planning.
@@
Building on the recognition of reproducibility challenges and unstable RL training of LLM agents, recent studies have proposed best practices for training LLM agents using RL methods. dang2025rlforreasoning recommend leveraging high quality data, balancing easy and hard problems, and controlling length generation with cosine reward. yu2025dapo promote higher clipping in the GRPO loss to promote diversity and avoid entropy collapse, dynamic sampling to improve training efficiency and stability, token level gradients for long CoT sequences, and overlong reward shaping to reduce reward noise. leroux2025tapered introduce tapered variant of importance sampling to speed up learning while maintaining stable learning dynamics. The proposed method (TOPR) allows the handling of both positive and negative examples in a fully offline setting. More generally, hochlehnert2025repro emphasizes the need for greater methodological precision, particularly concerning decoding parameters, random seeds, prompt formatting, as well as the hardware and software frameworks, to guarantee transparent and thorough assessments of model performance.
@@
Recent advancements have sought to bridge the gap in training LLM agents for multi-step environments, with approaches like WebRL qi2025webrl and SWEET-RL zhou2025sweetrltrainingmultiturnllm demonstrating significant progress. WebRL employs a self-evolving curriculum to address the challenges of sparse feedback and task scarcity, successfully enhancing the performance of open LLMs in web-based tasks qi2025webrl. Similarly, SWEET-RL introduces a hierarchical structure that enables effective credit assignment over multiple turns, improving policy learning and generalization in collaborative reasoning tasks zhou2025sweetrltrainingmultiturnllm. These studies collectively illustrate the necessity of adapting RL techniques to accommodate the complexities of multi-step interactions, paving the way for more capable and versatile LLM agents.